the shortest distance between line y-x=1 and the curve x=y2

Grade:Select Grade

1 Answers

Aziz Alam IIT Roorkee
8 years ago
Any point on parabola, (k2,k)
Perpendicular distance formula:
D=(k-k2-1)/21/2
Differentiating and putting =0
1-2k=0
k=1/2
Therefore the point is (1/4, 1/2)
D=3/(321/2)